# Categories of Line Defects and Cohomological Hall Algebras

### v2

@@ 1.1 Half-BPS lines and the Holomorphic-Topological twist
We expect half-BPS line defects to map to special topological line defects. Indeed, they preserve twice the amount of supersymmetry that is needed for a line defect to survive the twist. Although we do not know how to make this expectation concrete from first principles, we can take inspiration from a rigorous mathematical definition cautis2023canonical of a monoidal category $\mathrm{KP}_{G,N}$, which conjecturally describes the image of half-BPS line defects in the HT twist of Lagrangian SQFTs whose gauge group is the compact form of a reductive group $G$ and which has matter fields (hypermultiplets) valued in $T^{*}N$ for some representation $N$ of $G$.
This category sits in a bigger dg (triangulated) category as the heart of a t-structure. Roughly, this means that the Chan-Paton factors one encounters in the monoidal structure are vector spaces rather than complexes. [^2] When a calculation is done in the physical theory and involves half-BPS configurations, it will involve Hilbert spaces which are representations of the physical spin and of the R-charge $SU(2)_{R}$ symmetry algebra, as well as four super-charges. The twisted description only keeps track of the Cartan generator of spin, denoted as $n$, and the Cartan generator of $SU(2)_{R}$, which defines the ghost number. A no-exotics conjecture claims that the ground states which contribute to Chan-Paton factors, etc. carry trivial representations of $SU(2)_{R}$, in agreement with the mathematical perspective. The no-exotics conjecture is compatible with other conjectures present in this paper, but is still rather mysterious. It would be interesting to know how the mathematical formalism could be adjusted to deal with situations where the no-exotics conjecture might be false.
@@
Accordingly, we introduce an “algebra of BPS states” $A_{u}$, perhaps in the sense of Harvey:1995fq , exemplified by the Cohomological Hall Algebra Kontsevich:2008fj $H(Q,W)$ of a quiver $Q$ with superpotential $W$ [^5] The positive subalgebra of the quiver Yangian defined in Li:2020rij can be viewed as the equivariant version of the spherical CoHA, see (Galakhov:2021vbo, , Sec. 5.2). For general quiver and potential, the relation between the ordinary CoHA and the equivariant version is subtle; however, as we will see later, for the 2-acyclic quivers that are the focus of the current paper, one can simply set all the equivariant parameters in the equivariant spherical CoHA to zero to obtain the ordinary one, see Appendix A. This is the algebra of BPS states for a SQFT which arises from String Theory, where BPS particles are D-branes whose dynamics is encoded in $H(Q,W)$. In Appendix B we will attempt to give a more direct physical definition of $A_{u}$.
@@ 1.2 Future directions and missing links
We could not make this idea work and found instead the alternative route of a bimodule category. It would nevertheless be interesting to identify such a description for the category of line defects.
@@ 2.1 Quivers and Superpotentials
  • If $(Q,W)$ is infinitely mutable in the sense of Kontsevich:2008fj ; derksen2008quivers reviewed below, $H(Q,W)$ is actually independent of the choice of $W$. Furthermore, it coincides with the “spherical part” $\mathrm{sH}(Q)$ of $H(Q,0)$, which has a simple algebraic description.
Accordingly, whenever we refer to the “CoHA” in the rest of the paper, we are actually referring to $\mathrm{sH}(Q)$ as a conjectural stand-in for $H(Q,W)$ for an infinitely mutable $(Q,W)$.
@@ 2.2 Quivers and Mutations
The mutations may be obstructed if the superpotential $W$ misses certain terms. Reference derksen2008quivers proves that there exist sufficiently generic superpotentials $W$ such that any arbitrary sequence of mutations can be executed, as long as $Q$ has no loops of length $1$ and $2$. We call such a pair $(Q,W)$ infinitely mutable. Whenever $W$ is not mentioned, we assume that it is generic and thus infinitely mutable.
@@ 2.3 Stability conditions and walls
The stability conditions are controlled by the central charges $Z_{\gamma}(u)$, which define a linear map $\Gamma_{u}\to\mathbb{C}$. The central charges are locally holomorphic and non-vanishing on the non-singular part of ${\mathcal{B}}$. In the rest of the paper we will never really need to refer directly to $u$: only $\Gamma$ and $Z_{\gamma}$ enter our considerations. We will thus drop any reference to $u$ in the following.888 In general, not all values of the central charges occur in ${\mathcal{B}}$. Accordingly, not all possible quiver mutations will occur along ${\mathcal{B}}$ Cecotti:2011rv . The implications of this fact are unclear.
@@
if we assume the absence of cycles of length $1$ and $2$. Equivalently, $|i\to j|=\langle\gamma_{i},\gamma_{j}\rangle$ when the latter is positive, or 00 otherwise. The collection of $\gamma_{i}$ in $\Gamma$ thus gives us directly the BPS quiver $Q$, and we equip it with a generic infinitely-mutable superpotential. We will denote a BPS quiver $Q$ with specific $\gamma_{i}$ labels as a labelled quiver.
Next, we can describe how the BPS quiver mutates across a wall of marginal stability of the second kind. The only way a populated central charge can hit the real axis is if it is the central charge $Z_{\gamma_{i}}$ of an elementary particle. Across the wall, the anti-particle with charge $-\gamma_{i}$ becomes an “elementary” particle, and the roles of elementary particles and bound states get re-shuffled.
@@ 2.5 Harder-Narasimhan filtrations and factorization
Now, suppose we approach a wall of marginal stability of the second kind, where $Z_{\gamma_{i}}$ becomes real for a node $i$. Then the factor $S_{\gamma_{i}}$ appears either leftmost (when $Z_{\gamma_{i}}<0$) or rightmost (when $Z_{\gamma_{i}}>0$) in the CoHA factorization. The form of this factor is independent of the quiver: it essentially involves representations of a quiver with a single node and no arrow. It is an infinite exterior algebra with generators $e^{(i)}_{n}$ of spin $12n+$ 1 and charge $\gamma_{i}$:
$S_{\gamma_{i}}=\Lambda^{\bullet}[e^{(i)}_{0},e^{(i)}_{1},e^{(i)}_{2},\cdots]\,.$ (30)
@@
$\Phi(X_{\gamma_{i}})\equiv\prod_{m=0}^{\infty}(1+X_{\gamma_{i}}\mathfrak{q}^{2m+1})=\sum_{n=0}^{\infty}\frac{\mathfrak{q}^{n^{2}}}{(\mathfrak{q}^{2};\mathfrak{q}^{2})_{n}}X_{n\gamma_{i}}\,.$ (31)
Accordingly, before the wall, we have
${\mathcal{S}}=\begin{cases}\begin{aligned} &\Phi(X_{\gamma_{i}})\left[\prod^{\curvearrowright}_{\gamma>\gamma_{i}}\chi(S_{\gamma})\right]\qquad\qquad Z_{\gamma_{i}}<0\\ &\left[\prod^{\curvearrowright}_{\gamma<\gamma_{i}}\chi(S_{\gamma})\right]\Phi(X_{\gamma_{i}})\qquad\qquad Z_{\gamma_{i}}>0\,,\end{aligned}\end{cases}$ (32)
@@ 2.6 Spherical CoHA and Spherical Shuffle Algebra
  • The Spherical CoHa $\mathrm{sH}(Q)$ coincides with $H(Q,W)$ for infinitely mutable $W$.
@@
The spin grading of $[p]_{\vec{n}}$ receives three contributions:
@@
We also assign a Grassmann parity $\sum_{i}n_{i}$ to $[p]_{\vec{n}}$, consistent with our physical interpretation.
@@
$[p_{1}(s^{(1)})]_{\vec{n}^{(1)}}\cdot[p_{2}(s^{(2)})]_{\vec{n}^{(2)}}\equiv\left[\sum_{\vec{\sigma}}\epsilon(\vec{\sigma})\,(\mathrm{fac}\times p_{1}\times\,p_{2})(\vec{\sigma}(s))\right]_{\vec{n}^{(1)}+\vec{n}^{(2)}}\,,$ (39)
@@
where $\sim$ means up to regular terms. This is the $h_{I}\rightarrow 0$ limit of the $e-e$ relations from the corresponding quiver Yangians, see (427).
@@ 2.8 Extra symmetries
$[p]_{\vec{n}}\to\left[p\prod_{i}\prod_{a_{i}=1}^{n_{i}}f_{i}(s_{i,a_{i}})\right]_{\vec{n}}$ (53)
@@ 3.2 The $A_{3}$ quiver and its relatives
This three-node quiver and its mutation orbit $[A_{3}]$ describes the BPS spectrum of the next simplest AD theory and illustrates nicely the difference between $\mathrm{Sh}$ and $\mathrm{sH}$ in the presence of loops.
@@
    $[s_{2,1}^{k}]_{(0,1,1)}\qquad\textrm{and}\qquad[s_{3,1}^{k}]_{(0,1,1)}\,,$ (106)
@@ 4 Framed BPS states and CoHA bimodules from framed quivers
that are connected to the new node *. We also include a new superpotential $W_{\vec{f}}$, which we assume to be infinitely-mutable and can treat it as before.
@@ 4.1 Shuffle bimodules
A generic element of the bimodule is denoted as $[p]_{\vec{n}}$, with spin grading
@@
We also assign a Grassmann parity $\sum_{i}n_{i}$ to $[p]_{\vec{n}}$. Finally, we also use the notation
@@ 4.4 PBW bases, factorizations, and mutations
where $\gamma<\theta$ means $\textrm{arg}Z_{\gamma}>\theta$ and vice versa. Again, the graded component $V^{\vartheta}(Q;\vec{f})_{\vec{f},\bullet}$ consists of $b$ only. We expect $V^{\vartheta}(Q;\vec{f})_{\vec{f},\bullet}$ to also admit a definition in terms of “stable framed representations of $(Q,W)$Soibelman:2014gta , but we will not explore this idea here.
@@ 4.5 Tensor products and factorization
so we would only land outside $\mathrm{sH}(Q;\vec{f}+\vec{f}^{\prime})$ if $\pi_{\vec{f},\vec{f}^{\prime}}(\widetilde{v}\otimes\widetilde{v}^{\prime})$ lands outside.
@@ 4.6 IR gauge equivariance
The bimodules that we have defined in this section also enjoy the obvious action of the $\Gamma^{\vee}({\mathcal{O}})$ symmetry algebra, which commutes with the bimodule action. Our tensor product decompositions are also compatible with the $\Gamma^{\vee}({\mathcal{O}})$ action.
@@ 4.7 A test of invertibility for $\mathrm{RG}_{u}$
where “$\mathrm{Tr}$” denotes the projection $X_{\gamma}\to\delta_{\gamma,0}$.
@@ 5 Framed BPS degeneracies, factorizations, and tensor products for the $A_{1}$ quiver
The 4D Coulomb branch of the theory has two chambers. If we pick one of the two as a reference, denoting by $\gamma_{1}$ the gauge charge of the hypermultiplet with $\mathrm{Im}\,Z_{\gamma_{1}}>0$ in that chamber, the second chamber is reached by a mutation of the node and the elementary charge is $-\gamma_{1}$. The full charge lattice can be identified as $\Gamma=\mathbb{Z}^{2}$, with only the second factor being populated by BPS particles: $\gamma_{1}=(0,1)$. We define a pairing $\langle(a,b),(c,d)\rangle=ad-bc$.
@@
The algebra of equivariant K-theory classes $[L_{m,e}]$ is generated by the Wilson lines $[L_{0,n}]=[L_{0,1}]^{n}$ for integer $n$ and the ’t Hooft lines $[L_{1,0}]$ and $[L_{-1,0}]$, which satisfy

### v3

@@ 1.1 Half-BPS lines and the Holomorphic-Topological twist
We expect half-BPS line defects to map to special topological line defects. Indeed, they preserve twice the amount of supersymmetry that is needed for a line defect to survive the twist. Although we do not know how to make this expectation concrete from first principles, we can take inspiration from a rigorous mathematical definition cautis2023canonical of a monoidal category $\mathrm{KP}_{G,N}$, which conjecturally describes the image of half-BPS line defects in the HT twist of Lagrangian SQFTs whose gauge group is the compact form of a reductive group $G$ and which has matter fields (hypermultiplets) valued in $T^{*}N$ for some representation $N$ of GG. [^2] While a general prescription for associating a BPS quiver to an arbitrary $\mathrm{KP}_{G,N}$ is not yet known, many important cases can be handled explicitly. For instance, any pure gauge theory with a unitary gauge group has a known BPS quiver. More recently, Ambrosino:2025qpy gave a procedure to obtain the BPS quiver of any gauge theory with cotangent matter using the BPS quiver of the pure gauge theory as input. Together, these results cover a large and important sub-class of cautis2023canonical , and we expect them to extend further.
The category $\mathrm{KP}_{G,N}$ sits in a bigger dg (triangulated) category as the heart of a t-structure. Roughly, this means that the Chan-Paton factors one encounters in the monoidal structure are vector spaces rather than complexes. [^3] When a calculation is done in the physical theory and involves half-BPS configurations, it will involve Hilbert spaces which are representations of the physical spin and of the R-charge $SU(2)_{R}$ symmetry algebra, as well as four super-charges. The twisted description only keeps track of the Cartan generator of spin, denoted as $n$, and the Cartan generator of $SU(2)_{R}$, which defines the ghost number. A no-exotics conjecture claims that the ground states which contribute to Chan-Paton factors, etc. carry trivial representations of $SU(2)_{R}$, in agreement with the mathematical perspective. The no-exotics conjecture is compatible with other conjectures present in this paper, but is still rather mysterious. It would be interesting to know how the mathematical formalism could be adjusted to deal with situations where the no-exotics conjecture might be false.
@@
Accordingly, we introduce an “algebra of BPS states” $A_{u}$, perhaps in the sense of Harvey:1995fq , exemplified by the Cohomological Hall Algebra Kontsevich:2008fj $H(Q,W)$ of a quiver $Q$ with superpotential WW. [^6] The positive subalgebra of the quiver Yangian defined in Li:2020rij can be viewed as the equivariant version of the spherical CoHA, see (Galakhov:2021vbo, , Sec. 5.2). For general quiver and potential, the relation between the ordinary CoHA and the equivariant version is subtle; however, as we will see later, for the 2-acyclic quivers that are the focus of the current paper, one can simply set all the equivariant parameters in the equivariant spherical CoHA to zero to obtain the ordinary one, see Appendix A. This is the algebra of BPS states for a SQFT which arises from String Theory, where BPS particles are D-branes whose dynamics is encoded in $H(Q,W)$. In Appendix B we will attempt to give a more direct physical definition of $A_{u}$.
@@ 1.2 Future directions and missing links
We could not make this idea work and found instead the alternative route of a bimodule category. It would nevertheless be interesting to identify such a description for the category of line defects. Namely, there might be a construction of modules for the “double” of the CoHA or Yetter-Drinfeld modules of CoHA that could provide an alternative description of the category of line defects. As far as we know, there is currently no well-defined notion of such a double or of a “true” coproduct on CoHA suitable for our needs (despite much effort), particularly when one takes into account the additional Heisenberg grading. We can nevertheless speculate about the form this equivalence might take. Recall that for a Hopf algebra $A$, there is a monoidal equivalence between the category of (left–left) Yetter–Drinfeld modules over $A$ and the category of $A$-bicomodules in the category of $A$-bimodules; under this equivalence, an object $V$ is sent to $V\otimes A$. In particular, finite-dimensional $A$-modules correspond to bimodules that are free and finitely generated as right $A$-modules – precisely the form of the bimodules that appeared in our work. For CoHAs, the above correspondence would require a “chiral” version of all comodule structures. The existence and significance of such a chiral bicomodule structure on our CoHA-bimodules certainly deserve further study, and we hope to investigate this in the future.
@@ 2.1 Quivers and Superpotentials
  • For every quiver there exists at least one potential $W$ that is infinitely mutable (in the sense of Kontsevich:2008fj ; derksen2008quivers , see review below) for which the CoHA is spherically generated, namely $H(Q,W)$ coincides with the “spherical part” $\mathrm{sH}(Q)$ of $H(Q,0)$, which has a simple algebraic description.
Accordingly, whenever we refer to the “CoHA” in the rest of the paper, we are actually referring to $\mathrm{sH}(Q)$ as a conjectural stand-in for $H(Q,W)$ for the suitable infinitely mutable $(Q,W)$.
Note added for v3. In earlier versions of the paper, the conjecture was stated as “$H(Q,W)$ is spherically generated and coincides with the “spherical part” $\mathrm{sH}(Q)$ of $H(Q,0)$ whenever $W$ is infinitely-mutable. It was shown by Davison in Davison:2025ydh that for a given $Q$, this might not hold for every infinitely-mutable WW”. Take the example of the Markov quiver $Q$, although it admits an infinitely-mutable potential $W_{\mathrm{Type-4}}$ (in (Davison:2025ydh, , Lemma 2.1.)) for which $H(Q,W_{\mathrm{Type-4}})$ is spherical and $H(Q,W_{\mathrm{Type-4}})=\textrm{sH}(Q)$, it also admits another infinitely mutable superpotential $W_{\mathrm{Type-5}}$ (in (Davison:2025ydh, , Lemma 2.1.)) for which $H(Q,W_{\mathrm{Type-5}})$ strictly contains its spherical subalgebra.
Note that both Type 4 and Type 5 correspond to a 4d $\mathcal{N}=2$ theory: Type 4 to a “minimal” theory and Type 5 to the combination of the minimal theory and a decoupled sector. (This would also be a reason to prefer Type 4 to Type 5.) The corresponding algebraic statement is that $H(Q,W_{\mathrm{Type-5}})$ splits as the product of $\textrm{sH}(Q)$ and a free exterior algebra generated by the non-spherical elements.”
Accordingly, Davison in Davison:2025ydh has proposed to modified the conjecture into the current form. This refined conjecture is sufficient for our purposes (perhaps combined with the conjecture “any infinitely mutable potential can be related to a valid 4d N=2 theory”), since our goal is to use CoHAs to define category of line defects for “some” 4d $\mathcal{N}=2$ theories.
@@ 2.2 Quivers and Mutations
The mutations may be obstructed if the superpotential $W$ misses certain terms. Reference derksen2008quivers proves that there exist sufficiently generic superpotentials $W$ such that any arbitrary sequence of mutations can be executed, as long as $Q$ has no loops of length $1$ and $2$. We call such a pair $(Q,W)$ infinitely mutable. Whenever $W$ is not mentioned, we assume that it is infinitely mutable. [^9] Note that we do not need to demand the stronger (and vaguer) “genericity” condition on $W$. Moreover, based on the conjecture in Sec. 2.1, we choose the infinitely-mutable $W$ for which $H(Q,W)$ is spherical and coincide with the spherical subalgebra $\textrm{sH}(Q)$ of $H(Q,0)$.
@@ 2.3 Stability conditions and walls
The stability conditions are controlled by the central charges $Z_{\gamma}(u)$, which define a linear map $\Gamma_{u}\to\mathbb{C}$. The central charges are locally holomorphic and non-vanishing on the non-singular part of $\mathcal{B}$}. In the rest of the paper we will never really need to refer directly to $u$: only $\Gamma$ and $Z_{\gamma}$ enter our considerations. We will thus drop any reference to $u$ in the following. [^10] In general, not all values of the central charges occur in $\mathcal{B}$}. Accordingly, not all possible quiver mutations will occur along $\mathcal{B}$} Cecotti:2011rv . The implications of this fact are unclear.
@@
if we assume the absence of cycles of length $1$ and $2$. Equivalently, $|i\to j|=\langle\gamma_{i},\gamma_{j}\rangle$ when the latter is positive, or 0 otherwise. The collection of $\gamma_{i}$ in $\Gamma$ thus gives us directly the BPS quiver $Q$, and we equip it with an infinitely-mutable superpotential $W$ for which $H(Q,W)$ is spherical and coincide with the spherical subalgebra $\textrm{sH}(Q)$ of $H(Q,0)$. We will denote a BPS quiver $Q$ with specific $\gamma_{i}$ labels as a labelled quiver.
Next, we can describe how the BPS quiver mutates across a wall of marginal stability of the second kind. The only way a populated central charge can hit the real axis is if it is the central charge $Z_{\gamma_{i}}$ of an elementary particle. Across the wall, the anti-particle with charge $\gamma_{i}$ becomes an “elementary” particle, and the roles of elementary particles and bound states get re-shuffled.
@@ 2.5 Harder-Narasimhan filtrations and factorization
Now, suppose we approach a wall of marginal stability of the second kind, where $Z_{\gamma_{i}}$ becomes real for a node $i$. Then the factor $S_{\gamma_{i}}$ appears either leftmost (when $Z_{\gamma_{i}}<0$) or rightmost (when $Z_{\gamma_{i}}>0$) in the CoHA factorization. The form of this factor is independent of the quiver: it essentially involves representations of a quiver with a single node and no arrow. It is an infinite exterior algebra with generators $e^{(i)}_{n}$ of spin $2n+1$ and charge $\gamma_{i}$:
$S_{\gamma_{i}}=\Lambda^{\bullet}[e^{(i)}_{0},e^{(i)}_{1},e^{(i)}_{2}$$\cdots$]\,. (30)
@@
$\Phi(X_{\gamma_{i}})\equiv\prod_{m=0}^{\infty}(1+X_{\gamma_{i}}\mathfrak{q}^{2m+1})=\sum_{n=0}^{\infty}\frac{\mathfrak{q}^{n^{2}}}{(\mathfrak{q}^{2};\mathfrak{q}^{2})_{n}}X_{n\gamma_{i}}\,$, (31)
where $(\mathfrak{q}^{2};\mathfrak{q}^{2})_{n}=\prod^{n}_{k=1}(1-\mathfrak{q}^{2k})$. Accordingly, before the wall, we have
[…]
@@ 2.6 Spherical CoHA and Spherical Shuffle Algebra
  • The Spherical CoHa $\mathrm{sH}(Q)$ coincides with $H(Q,W)$ for some infinitely mutable $W$.
@@
The spin grading of $p]_{\vec{n}}$ receives three contributions:
@@
We also assign a Grassmann parity $\sum_{i}n_{i}$ to $p]_{\vec{n}}$, consistent with our physical interpretation.
@@
$p_{1}(s^{(1)})]_{\vec{n}^{(1)}}\cdot[p_{2}(s^{(2)})]_{\vec{n}^{(2)}}\equiv\left[\sum_{\vec{\sigma}}\epsilon(\vec{\sigma})\,(\mathrm{fac}\times p_{1}\times\,p_{2})(\vec{\sigma}(s))\right]_{\vec{n}^{(1)}+\vec{n}^{(2)}}\,$, (39)
@@
where $\sim$ means up to regular terms. This is the $h_{I}\rightarrow 0$ limit of the eee-e relations from the corresponding quiver Yangians, see (427).
@@ 2.8 Extra symmetries
$p]_{\vec{n}}\to\left[p\prod_{i}\prod_{a_{i}=1}^{n_{i}}f_{i}(s_{i,a_{i}})\right]_{\vec{n}}$ (53)
@@ 3.2 The $A_{3}$ quiver and its relatives
This three-node quiver and its mutation orbit $A_{3}]$ describes the BPS spectrum of the next simplest AD theory and illustrates nicely the difference between $\mathrm{Sh}$ and $\mathrm{sH}$ in the presence of loops.
@@
    $s_{2,1}^{k}]_{(0,1,1)}\qquad\textrm{and}\qquad[s_{3,1}^{k}]_{(0,1,1)}\,$, (106)
@@ 4 Framed BPS states and CoHA bimodules from framed quivers
that are connected to the new node *. We also include a new superpotential $W_{\vec{f}}$, which we assume to be infinitely-mutable and is such that $H(Q_{\vec{f}},W_{\vec{f}})$ is spherical and is isomorphic to the spherical subalgebra $\textrm{sH}(Q_{\vec{f}})$ of $H(Q_{\vec{f}},0)$, and can thus be treated as before.
@@ 4.1 Shuffle bimodules
A generic element of the bimodule is denoted as $p]_{\vec{n}}$, with spin grading
@@
We also assign a Grassmann parity $\sum_{i}n_{i}$ to $p]_{\vec{n}}$. Finally, we also use the notation
@@ 4.4 PBW bases, factorizations, and mutations
where $\gamma<\theta$ means $\textrm{arg}Z_{\gamma}>\theta$ and vice versa. Again, the graded component $V^{\vartheta}(Q;\vec{f})_{\vec{f},\bullet}$ consists of $b$ only. We expect $V^{\vartheta}(Q;\vec{f})_{\vec{f},\bullet}$ to also admit a definition in terms of “stable framed representations of (Q,W)(Q,W)Soibelman:2014gta , but we will not explore this idea here.
@@ 4.5 Tensor products and factorization
[…]
@@ 4.6 IR gauge equivariance
The bimodules that we have defined in this section also enjoy the obvious action of the $\Gamma^{\vee}({\mathcal{O}})$ symmetry algebra (defined in Sec. 2.8), which commutes with the bimodule action. Our tensor product decompositions are also compatible with the $\Gamma^{\vee}({\mathcal{O}})$ action.
@@ 4.7 A test of invertibility for $\mathrm{RG}_{u}$
where $\mathrm{Tr}$” denotes the projection $X_{\gamma}\to\delta_{\gamma,0}$, and the prefactor $(\mathfrak{q}^{2})^{\mathrm{rk}(\Gamma)}_{\infty}$ comes from the contribution from the ghosts $c_{i,n}$ with $i=1,\dots,\mathrm{rk}(\Gamma)$.
@@ 5 Framed BPS degeneracies, factorizations, and tensor products for the $A_{1}$ quiver
The 4D Coulomb branch of the theory has two chambers. If we pick one of the two as a reference, denoting by $\gamma_{1}$ the gauge charge of the hypermultiplet with $\mathrm{Im}\,Z_{\gamma_{1}}>0$ in that chamber, the second chamber is reached by a mutation of the node and the elementary charge is $\gamma_{1}$. The full charge lattice can be identified as $\Gamma=\mathbb{Z}^{2}$, with only the second factor being populated by BPS particles: $\gamma_{1}=(0,1)$. We define a pairing $\langle(a,b),(c,d)\rangle=ad-bc$.
@@
The algebra of equivariant K-theory classes $L_{m,e}]$ is generated by the Wilson lines $L_{0,n}]=[L_{0,1}]^{n}$ for integer $n$ and the ’t Hooft lines $L_{1,0}]$ and $L_{-1,0}]$, which satisfy